OpenStudy (anonymous):

Find an equation of a sphere whose diameter has endpoints with coordinates (1,0,3) and (3,-5,6)

OpenStudy (anonymous):

what's up dude

OpenStudy (anonymous):

hey, what's this?

OpenStudy (anonymous):

reviewing for my final

OpenStudy (anonymous):

this was a question from the study guide i was never able to solve

OpenStudy (anonymous):

find center point between them (1,0,3) and (3,-5,6) center of sphere (2,-5/2,9/2)

OpenStudy (anonymous):

(x-2)^2 + (y+5/2)^2+(z-9/2)^2=r^2

OpenStudy (anonymous):

shouldn't i use the distance formula to get the radius?

OpenStudy (anonymous):

for radius^2, (2-1)^2 + (-5/2)^2 +(9/2)^2

OpenStudy (anonymous):

1+25/4 +81/4

OpenStudy (anonymous):

(x-2)^2 + (y+5/2)^2+(z-9/2)^2=1+25/4 +81/4

OpenStudy (anonymous):

so 96/4?

OpenStudy (anonymous):

is r?

OpenStudy (anonymous):

r^2

OpenStudy (anonymous):

sphere formula x^2 +y^2+z^2=r^2

OpenStudy (anonymous):

so my equation would be in the form x^2 + y^2 + z^2 = (96/4)^2?

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 + (z-a)^2 = (110/4)^2? 4+25+81=110

OpenStudy (anonymous):

lol oops

OpenStudy (anonymous):

alright...so how did you get ur center?

OpenStudy (anonymous):

it looks like you subrtacted the x values....but then added y and z values and divided by 2

OpenStudy (anonymous):

by finding center between two endpoints of diameter

OpenStudy (amistre64):

midpoint; add points and divide by 2

OpenStudy (anonymous):

ohhhhh lol

OpenStudy (anonymous):

i see...didn't think that through clearly

OpenStudy (anonymous):

(1,0,3) and (3,-5,6) (1+3 )/2, (0-5)/2, (3+6)/2)

OpenStudy (amistre64):

i go for the vertical format meself :) (1, 0, 3) (3,-5,6) ------- (4,-5,9) / 2